OpenStudy (anonymous):

Which of the following demonstrates the Distributive Property? 2(3a + 4) = 6a + 4 2(3a + 4) = 3a + 8 2(3a + 4) = 6a + 8 2(3a + 4) = 5a + 6

OpenStudy (welshfella):

the 2 multiplies each term in the parentheses

OpenStudy (anonymous):

In general the distributive property says\[a \left( b+c \right)=ab+ac\] Choose the answer that shows this property.